Beth, Kinsley, and Kori went shopping. Beth spent half of her money on a new pair of khaki pants, then she gave Kinsley $4.00 that she owed her. Beth spent 1/4 of her remaining money on lunch. After lunch, she spent 2/3 of her remaining money on a tank top. Finally, she spends $3.00 on popcorn and a soda. She went home with $4.00. How much money did she start the day with?
Answer by mananth(16949)

let Beth have $x
she spent x/2 on khaki pants.
she had balance x/2.
....
she gave Kinsley $4
x/2 - 4 is the balance she had now.
..
1/4((x/2-4) she spent on lunch.
so she had a balance of
(x/2)-4 - (1/4(x/2 -4)
..
2/3 [(x/2)-4 - (1/4(x/2 -4)] she spent on tank top
and spent $3 for popcorn and soda
2/3 [(x/2)-4 - (1/4(x/2 -4)]-3 =4
2/3 [(x/2)-4 - (1/4(x-8)/2]-3 =4
2/3 [(x/2)-4 - ((x-8)/8]-3 =4
LCD =8
2/3 [(4x-32 -x+8)/8]-3 =4
2/3 [(3x-24)/8]-3 =4
[(3x-24)/12] -3=4
[(3x-24)/12]=4+3
[(3x-24)]=7*12
3x-24=84
3x=108
x= 36
She had $36 before shopping